# Detecting and Mitigating Bias in Algorithms Used to Disseminate Information in Social Networks

### v1

@@
Social connections are a conduit through which individuals communicate, information propagates, and diseases spread. Identifying individuals that are more likely to adopt ideas or technologies and spread them to others is essential in order to develop effective information campaigns, fight epidemics, and to maximize the reach of limited resources. Consequently a lot of work has focused on identifying sets of influencers. Here we show that seeding information using these influence maximization methods, only benefits connected and central individuals, consistently leaving the most vulnerable behind. Our results highlights troublesome outcomes of influence maximization algorithms: they do not disseminate information in an equitable manner threatening to create an increasingly unequal society. To overcome this issue we devise a simple, multi-objective algorithm, which maximises both influence and information equity. Our work demonstrates how to find fairer influencer sets, highlighting that in our search for maximizing information, we do not need to compromise on information equality.
@@ Introduction
Social relationships serve as important vectors through which a multitude of behaviors spread, from health related behaviors  (?, ?), innovation  (?, ?), decisions of micro-financing  (?), decision to insure  (?), happiness  (?), cultural tastes  (?), to the emergence of social movements  (?, ?). Knowing through which pathways information spreads is vital for international development  (?, ?) and crucial in developing more efficient methodologies that maximize the diffusion of potentially life saving information  (?, ?). Due to resource constraints it is unfeasible to send a piece of information to all individuals within a network. Instead, a frequently adopted strategy is to seed a small set of individuals, much smaller than the full population, located at strategic places in the network whose activation (or removal) would facilitate the spread of information (or in the case of epidemics inhibit a disease from spreading). Numerous methods have been proposed to identify ”the set” of influential nodes. These methodologies can be divided up into two fundamentally distinct classes  (?, ?): (1) methods that identify individuals who are highly connected or central, and effective at diffusing information – so-called superspreader methods  (?, ?, ?, ?, ?, ?, ?). And (2) methods which highlight individuals that occupy structurally vital positions in a network whose removal would destroy the network and subsequently block information from propagating – also called superblocker methods  (?, ?, ?, ?, ?, ?). Although the methods are different there is a general consensus that they both pinpoint nodes which are highly efficient conduits for information propagation  (?).
@@
Unfortunately, we have a limited understanding of which demographics these methods reach and which communities they leave behind, but there are alarming signs in the network science literature. Previous work has demonstrated that wealth is connected to diversity of social connections, the richer you are the more diverse is your social network, in turn enabling you to receive information from many directions  (?, ?, ?). Similarly, influence maximisation has been shown create gaps in information access  (?), and to have a gender skew, with male individuals having an advantage in being selected as influencers  (?, ?, ?). Previous work has also shown that an individuals’ chances of being ranked as an influencer are highly correlated with personal economic status  (?). Further, as social systems display strong levels of homophily  (?), where connections between similar individuals occur at higher rates than between dissimilar individuals, wealthy individuals are more likely to befriend people that resemble them, and less likely to interact with the poor  (?). As a consequence, information in social networks tends to be localized within social strata, restricting diffusion across demographic and socioeconomic gaps and between majority and minority groups  (?). Access to information is a major factor of social vulnerability  (?). By naively using current influence maximisation methodologies to selects predominantly wealthy influencers, through which information is inserted, we run the risk of tailoring information campaigns towards the most affluent groups of our societies, while severely under-representing the most vulnerable and marginalized.
@@
Vulnerability is a complex issue determined by physical, economic, social, and environmental factors, which decrease the capacity of individuals and groups to cope, anticipate, and react to hazards  (?). Here we study one aspect of vulnerability, namely individuals’ access to information. Previous work has looked into the interplay between influence maximization and different characteristics of the nodes (such as age, gender, ethnicity) which receive information  (?, ?, ?, ?). However, rather than assuming we have access to this type of information we focus on the general utility of the information individuals receive—this has previously been called a ’welfare approach’  (?, ?). As such, we study how access to information is distributed across all nodes in the network. Specifically, we define informational vulnerability as the average likelihood that an individual receives information, estimated from many independent diffusion cascades. We focus on two facets of this stochastic process: 1) frequency: how often does an individual receive information (i.e. how often are they reached by cascades), and 2) recency: how old are the cascades when they reach them (e.g. at what step of the cascade is the individual reached by it). Looking at just one cascade will reveal very little, but upon many independent simulations it is easy to see that not every individual is equally lucky on their network position and that some tend to be reached by more cascades, or earlier than others. Network structure and node locations in the network play very important roles in this. To simulate information diffusion processes on networks we apply the commonly used Independent Cascade Model (ICM) in its most simple form, unweighted and undirected  (?). ICMs are commonly used to study influence maximisation in social networks  (?, ?) and are a special case of susceptible-infected-recovered (SIR) models, where the recovery probability is fixed to 1. The ICM allows an informed individual one attempt to speak to their neighbors and convince them to adopt a behavior; the neighbors, if convinced, will then try to convince their neighbors, and so on (see Supplementary Materials (SM) Sec. S3 for more details).
To identify sets of influencers we focus on four widely used methods: (i) highest degree  (?) (HD), which is a commonly used heuristic where nodes are selected according to their number of connections, (ii) degree discount  (?) (DD), an efficient method for identifying superspreaders, (iii) coreHD  (?) (CHD), a state-of-the-art method for inferring superblockers, and (iv) k-core  (?) (KC), selecting nodes located in the core of the network. Fig. 1a illustrates, for a small real-world network, which nodes each method pinpoints as influencers. As ICM is a stochastic process we average over multiple simulations. For each realisation of the dynamic process we track which nodes are activated and how long it takes for the spreading process to reach them. We quantify this using two measures. The first measure, information frequency:
@@
summarizes the average fraction of times node $i$ has been reached, where $a_{i,n}=1$ if node $i$ received information in realisation $n$ and is zero otherwise, and $M$ is the total number of simulations. Information frequency ($\nu_{i}$) lies in the interval 0 to 1, where zero indicates that a node is never reached by the $M$ cascades, while a value of 1 indicates the node is reached by every cascade.
@@
quantifies the temporal delay from process initialization ($t=0$) until node $i$ is activated. Recency is calculated as the average of the inverse activation time to handle cases where the information spreading process dies out before reaching a node. Nodes that on average receive information quickly have $t_{i,n}\rightarrow 0$ and $\tau\rightarrow 1$, while nodes that are activated very late, or never, ($t_{i,n}\rightarrow\infty$) have $\tau\rightarrow 0$.
@@
To understand the shortcomings of the four methods in selecting influencers we compare them to a benchmark model where all nodes have equal chance of being selected (selected at random) — we call this the effective measure. If the ratio $\nu_{i}^{\text{method}}/\nu_{i}^{\text{benchmark}}>1$ a node will on average receive information more frequently when seeds are selected using a specific method as compared to when nodes are selected at random. If $\nu_{i}^{\text{method}}/\nu_{i}^{\text{benchmark}}<1$ a node will be better off when information is inserted at random entry points in the network. The same holds for recency, $\tau_{i}^{\text{method}}/\tau_{i}^{\text{benchmark}}>1$ indicates that a node on average receives more recent information when using a influencer maximisation method, while $\tau_{i}^{\text{method}}/\tau_{i}^{\text{benchmark}}<1$ denotes that information is received faster when information is inserted at random nodes. Fig. 1b illustrates the resulting effective recency-values from using influencers inferred by the four methods as seeds. Independent on methodology, the influencer nodes and their surrounding neighbors are always reached by the influencer set, however, a large fraction of nodes seems to be left behind. Typically nodes located on the periphery.
@@
Figure 1: Inequalities and diffusion of information in networks. a, Initial seed sets selected according to HD, CHD, DD, and KC, illustrating variations in how the four methods select influencers for a social network between households in a south-Indian village  (?). In this example $5\%$ of nodes are selected as influencers. Colored nodes indicate selected influencers. b, Effective recency for the social network. Recency is estimated across 1000 runs with $p_{c}=0.069$. (se SM Sec. S3.1) c, Cumulative distribution of individual node frequency within synthetic scale-free (SF) networks with $N=10^{4}$, $\gamma=2.5$, and $\langle p_{c}\rangle=0.085$. The curves show the probability that $\nu$ is less than or equal to $x$, where $x$ is any arbitrary value. Results are combined over 100 different network realizations. For each network we select an initial influencer set (1 % of nodes) inferred from one of the heuristics, run the spreading process, track which nodes receive information, and repeat the process $M=10N$ times to account for stochasticity in both the seed selection and the spreading processes. Red shaded regions denote parts of the distribution where the effective measure (the ratio) is below 1, while grey shaded indicate places where the ratio is above 1. d, Cumulative distribution of recency within SF networks. e, Fraction of nodes that are worse off with respect to information frequency in $n$ of the seeding heuristics when compared to the benchmark. Error bars are standard deviation over 100 network realizations. f, Fraction of nodes that are worse off with respect to recency.
@@
To formalize the observations from Fig. 1b we first investigate the four methodologies on a testbed of 100 realizations of synthetic unweighted and undirected networks with scale-free (SF) degree distributions (see SM Sec. S6 for random networks with normal degree distributions). While perfect SF networks are rarely observed in nature they are powerful simplifications of real-world networks  (?, ?). In order to compare heuristics we construct influencer sets from a fixed finite fraction of the network population—1% of nodes (see SM Sec. S5 for other seed sizes).
@@
Seeding information through random nodes in SF networks results in a near-homogeneous frequency distribution: i.e. node receives a constant stream of information (Fig. 1c, black line). In Fig. 1c, a completely equal information metric would be characterized by a vertical line in the cumulative probability distribution. Note that the black line of Fig. 1c, characterizing the random process, is not vertical due to the intrinsic variations of the network structure. Using the four methods to select influencers sets, however, results in fundamentally different frequency distributions (Fig. 1c, colored lines). Approximately half of nodes have effective frequency values ($\nu^{\text{method}}/\nu^{\text{benchmark}}$) above one, meaning they receive information more frequently then expected using a random process, while half the network receives much less compared to when seeds are selected at random (see also SM Sec. S4). The situation is similar for the recency distribution, albeit slightly more polarized (Fig. 1d). Looking across the different influencer maximisation heuristics, Fig. 1e, demonstrates that if a node is under-informed by one method, it will most likely not be better reached by any other method. On average $42.8\pm 1.7\%$ of nodes are better informed when information is seeded using any influencer maximization method. We say that these nodes are always better off. However, $47.0\pm 2.2\%$ of nodes are consistently left behind, independent of which method is used to select influencers. With respect to recency, Fig. 1f highlights an even worse situation, $55.6\pm 2.2\%$ of nodes receive out-of-date information, independent on which information maximisation methodology is used.
@@
To understand the implications of information inequalities for real-world networks we look at communication, interaction, and social networks varying in size from hundreds to tens of thousands of individuals. The networks differ in context from: face-to-face encounters  (?), connections between bloggers and blogs on political topics  (?), digital communication between university students  (?), online friendships  (?), and scientific collaborations  (?) (see SM Sec. S2 for more details).
@@
For information frequency, Fig. 2a shows that current state-of-the-art influencer heuristics identify seed-sets that, on average, leave significant portions of the network in disadvantaged positions. HD, CHD, and DD methods result in fairly similar $\nu$ distributions, while selecting influencers according to KC performs much worse (approx. 80% of nodes are better reached by randomly selected influencers). This is due to real-world network having a large numbers of cliques, where small discrepancies in shell numbers can result in KC only selecting nodes from a single clique  (?), effectively limiting the diffusion of information. Similar behavior is observed for information recency ($\tau$). Fig. 2b shows that HD, CHD, and DD leave behind a comparable numbers of nodes, while KC again performs worse. Summarizing the average reach of influencer heuristics, we find that up to $69.8\%$ of a network might receive information less frequently compared to if it is input at random (Fig. 2c), and information can reach up to $79.2\%$ of individuals slower (Fig. 2d).
@@
There is a connection between access to information and a node’s position in the network  (?). The connection is so strong that a predictive model, based on structural features such as node degree, k-shell number, clustering, centrality, and eccentricity, can accurately predict whether a node will fall into the group of ‘worse off according to all influencer heuristics‘ or ‘better off in all’ (see SM Sec. S7). We can, on average, with 97.4% accuracy predict the information status of nodes regarding frequency, and with an average accuracy of 96.9% for recency. This demonstrates that current influencer heuristics suffer from biases which disadvantage low-connected and peripheral nodes.
@@
Figure 2: Inequalities and diffusion of information in real-world networks. a, Cumulative distribution of individual node frequency within five real-world networks, ordered according to size (number of nodes). Initial seeds sets comprise $1\%$ of the network, and information frequency is averaged over $10N$ runs (see SM Table S1 for $N$-values). Curves show the probability that $\nu$ is less than or equal to x, where x is any arbitrary value. b, Cumulative distribution of recency within the empirical networks. c, Fraction of nodes that are worse off with respect to frequency in $n$ of the seeding heuristics when compared to the a random seeding procedure ($\nu^{\text{method}}/\nu^{\text{benchmark}}$). The fraction of disadvantaged individuals are: face-to-face ($20.7\%$), political blogs ($17.8\%$), student communication ($25.1\%$), online friendships ($69.8\%$), and academic collaborations ($56.0\%$). d, Fraction of nodes that are worse off with respect to recency ($\tau^{\text{method}}/\tau^{\text{benchmark}}$) for $n$ methods. Disadvantaged percentages include: face-to-face ($48.3\%$), political blogs ($34.5\%$), student communication ($34.3\%$), online friendships ($79.2\%$), and academic collaborations ($68.7\%$).
Fair Influence Maximisation
@@
To bridge the information gap different strategies can be employed. Previous studies have shown that instead of using influencer algorithms to identify $s$ individuals, one can select slightly more individuals $s+x$, but at random  (?). Even for small $x$-values this has been shown to result in larger cascades. Another solution is to apply acquaintance methods (also called friendship-nomination, which work by selecting a random neighbor of a randomly selected node), which have been used to select influencers for maximising mass drug administration campaigns  (?), to seed information about maternal and child health  (?), and for inferring centrally located individuals suitable as monitors for detecting large scale disease outbreaks  (?). Lastly, maximization of social welfare objective functions, along with the implementation of different algorithms, has been proposed to bridge the information gap  (?).
@@
We embark on a different solution. Traditionally, influence maximisation has only focused on maximising a single objective function, information spread. However, as literature from the field of Artificial Intelligence shows, focusing solely on optimizing one parameter can lead to troubling and unfair outcomes  (?, ?, ?). Instead, we propose a multi-objective formulation of fair influence maximisation, where both spread and fairness are taken into account in the fitness of a candidate solution (i.e. the set of selected influencers). We measure fairness as the fraction of nodes that receive information at a higher frequency and speed than what is expected from the benchmark model, where influencers are selected at random. We say a node is vulnerable if it receive less information than what can be expected at random (i.e. $\nu_{i}^{\text{method}}/\nu_{i}^{\text{benchmark}}<1$). In the following we focus primarily on information frequency, but similar results can be obtained for recency. As such, the more fair an influencer set is, the fewer nodes will be vulnerable, so we maximise the number of non-vulnerable nodes. Analytically calculating how information will spread from a set of nodes according to the ICM model is a computationally hard problem (NP-hard)  (?). As such, we use an approximation of the fitness of a influencer set (see SM Sec. S8). To find fairer seeds we then use a genetic algorithm to solve the optimization problem (see Methods).
@@
Fig. 3a,b show an example for one specific network, however, our fair influence maximization method works equally well for other networks (Fig. 3e-h). For other networks (Fig. 3e-h) the traditional heuristics identify influencer sets which optimize cascade size, however, we find that these seed sets are not fair in terms of information equality (for numerical results see Supplementary Materials). It is possible to improve this, with large gains in fairness being achievable by trading off a small reduction in cascade size. Overall we find that a $1\%$ trade-off in reach can result in a decrease in number of vulnerable individual by: $1.9\%$ for the network of political blogs, $3.1\%$ for student communication, $10.5\%$ for online friendships, and up to $24\%$ in collaboration networks (here we have disregard the face-to-face network due to the low number of identified seed sets which makes it difficult to calculate the proper trade-off value).
Lastly, for the collaboration network (Fig. 3h), we find that using traditional heuristics are not close to the Pareto frontier. Meaning they are sub-optimal in term of both cascades sizes and fair information access. In this situation our algorithm can also be used to identify, previously undiscovered, seed sets which optimize cascade-size.
Figure 3: Fair influence maximization for a social network between households in a South Indian village. a, Theoretical Pareto front of optimal influencer sets found by our algorithm, compared to the the ones found by the influencer heuristics. We disregard KC as it consistently performs worse, both in terms of fairness and reach compared to the other heuristics. Higher values of non-vulnerable nodes indicate higher values of fairness. b, Numerical evaluation of influencer sets using ICM. Error bars are given as the standard deviation from 10 realization of $10N$ ICM simulations. c, Edge activations for the set of influencers identified by HD. Edges are colored and sized according to how often they were activated during $10N$ simulations. Nodes colored black are initial seed nodes. d, Edge activations for one of the fairer seed sets identified by our algorithm. A comprehensive comparison between the two seed sets if available in the SM Sec. S9.

### v2

@@
Social connections are conduits through which individuals communicate, information propagates, and diseases spread. Identifying individuals who are more likely to adopt ideas and spread them is essential in order to develop effective information campaigns, maximize the reach of resources, and fight epidemics. Influence maximization algorithms are used to identify sets of influencers. Based on extensive computer simulations on synthetic and ten diverse real-world social networks we show that seeding information using these methods creates information gaps. Our results show that these algorithms select influencers who do not disseminate information equitably, threatening to create an increasingly unequal society. To overcome this issue we devise a multi-objective algorithm which maximizes influence and information equity. Our results demonstrate it is possible to reduce vulnerability at a relatively low trade-off with respect to spread. This highlights that in our search for maximizing information we do not need to compromise on information equality.
@@ Introduction
Social relationships serve as important vectors through which a multitude of behaviors spread, from health related behaviors  (?, ?), innovation  (?), decisions of micro-financing  (?), happiness  (?), cultural tastes  (?), to the emergence of social movements  (?). Knowing through which pathways information spreads is vital for international development  (?) and crucial for developing efficient methodologies that maximize the diffusion of potentially life saving information  (?, ?). Due to resource constraints it is unfeasible to send a piece of information to all individuals within a network. Instead, a frequently adopted strategy is to seed a small set of individuals, much smaller than the full population, located at strategic places in the network whose activation (or removal) would facilitate the spread of information (or in the case of epidemics inhibit a disease from spreading). Numerous methods have been proposed to identify “this set of influential nodes”. The methodologies can be divided up into two fundamentally distinct classes  (?, ?), superspreader and superblocker methods. Superspreader methods identify individuals who are highly connected and effective at diffusing information  (?, ?, ?, ?, ?, ?). Superblocker methods identify individuals that occupy structurally vital positions in a network whose removal would destroy the network and subsequently block information from propagating  (?, ?, ?, ?, ?). Although the methods are different there is a general consensus that they both pinpoint nodes which are highly efficient conduits for information propagation  (?).
@@
Unfortunately, we have a limited understanding of which demographics these methods reach and which communities they leave behind, but there are alarming signs in the literature. Previous work has shown that an individuals’ chances of being ranked as an influencer are highly correlated with personal economic status  (?). Similarly, influence maximization methods has been shown to create gaps in information access  (?), and to have a gender skew with male individuals having an advantage in being selected as influencers  (?, ?, ?). Further, social systems display strong levels of homophily, where connections between similar individuals occur at higher rates than between dissimilar individuals, with individuals being more likely to befriend people that resemble them  (?). As a consequence, information in social networks tends to be localized within social strata, restricting diffusion across demographic and socioeconomic gaps. In general, access to information is a major factor of social vulnerability  (?). By naively using current influence maximization methodologies to select influencers we run the risk of tailoring information campaigns towards the most affluent groups of our societies, while under-representing the most vulnerable and marginalized.
@@
Vulnerability is a complex issue determined by physical, economic, social, and environmental factors, which decrease the capacity of individuals and groups to cope, anticipate, and react to hazards  (?). Here we study one aspect of vulnerability, namely individuals’ access to information. Previous work has looked into the interplay between influence maximization and characteristics of nodes (such as age, gender, ethnicity) which receive information  (?, ?, ?, ?). Here we adopt a different approach. Rather than assuming we have access to demographic information of nodes we focus on the general utility of the information individuals receive—this is also called a ‘welfare approach’  (?, ?). We study how access to information is distributed across all nodes in the network using numerical simulations. Specifically, we define informational vulnerability as the average likelihood that an individual receives information, estimated from many independent diffusion cascades. We focus on two aspects of this stochastic process: 1) frequency: how often does an individual receive information (i.e. how often are they reached by cascades), and 2) recency: how old are the cascades when they reach them (i.e. at what step of a cascade is an individual reached). To simulate information diffusion processes in social networks we apply the commonly used Independent Cascade Model (ICM) in its most simple form, unweighted and undirected. ICMs are commonly used to study influence maximization in social networks  (?, ?). The ICM allows an informed individual one attempt to convince their neighbors to adopt a behavior (according to some probability); if successful the neighbors will try to convince their neighbors, etc. (see Supplementary Materials (SM) Sec. S3).
To identify sets of influencers we focus on two state-of-the-art influence maximization methods: degree discount  (?) (DD) which is an efficient method for identifying superspreaders, and coreHD  (?) (CHD) which infers superblockers. For comparison purposes we also include two commonly used heuristic for selecting influencers: highest degree  (?) (HD), which selects nodes according to their number of connections, and k-core  (?) (KC) which selects nodes located in the core of the network. Fig. 1a shows, for a small real-world network, which nodes each method selects as influencers. As ICMs are stochastic we average over multiple realizations. For each realization of the dynamic process we track which nodes are activated and how long it takes for the spreading process to reach them. We quantify this using two measures. The first measure, information frequency:
@@
summarizes the average fraction of times node $i$ has been reached, where $a_{i,n}=1$ if node $i$ received information in realization $n$ and zero otherwise, and $M$ is the total number of realizations. Information frequency ($\nu_{i}$) lies in the interval 0 to 1, where zero indicates that a node is never reached by any cascade, while a value of 1 indicates the node is reached by all cascades.
@@
quantifies the temporal delay from process initialization ($t=0$) until node $i$ is reached. Recency is calculated as the average of the inverse activation time to handle cases where the information spreading process dies out before reaching a node. Nodes that on average receive information fast have $t_{i,n}\rightarrow 0$ and $\tau\rightarrow 1$, while nodes that are reached very late, or never, ($t_{i,n}\rightarrow\infty$) have $\tau\rightarrow 0$.
@@
To uncover the shortcomings of the four influencer heuristics we compare them to a benchmark model where all nodes have an equal chance of being selected (selected at random) — we call this the effective measure. If the ratio $\nu_{i}^{\text{method}}/\nu_{i}^{\text{benchmark}}>1$ node $i$ will on average receive information more frequently when seeds are selected using a specific influencer mazimization heuristic as compared to when nodes are selected at random. If $\nu_{i}^{\text{method}}/\nu_{i}^{\text{benchmark}}<1$ node $i$ will be better off when information is inserted at random entry points in the network. The same holds for recency, $\tau_{i}^{\text{method}}/\tau_{i}^{\text{benchmark}}>1$ indicates node $i$, on average, receives more recent information when using a influencer maximization method, while $\tau_{i}^{\text{method}}/\tau_{i}^{\text{benchmark}}<1$ denotes that information is received faster when information is inserted at random nodes. Fig. 1b illustrates the resulting effective recency-values from using influencers inferred by the four heuristics as seeds. Independent of influence maximization methodology, the influencer nodes and their surrounding neighbors are always reached by the influencer set, however, a large fraction of nodes seems to be left behind. Typically nodes located on the periphery.
@@
Figure 1: Information is unequally distributed in networks. a, Initial seed sets selected according to HD, CHD, DD, and KC, showing variations in how the four methods select influencers for a social network between households in a south-Indian village  (?). Here, $5\%$ of nodes (colored) are selected as influencers for illustrative purposes (1% otherwise). b, Effective recency for the social network. Recency is estimated across 1000 runs with infection probability $p=0.069$ (see SM Sec. S3.1). c, Cumulative distribution of information frequency for synthetic SF-networks with $N=10^{4}$, $\gamma=2.5$, and average infection probability $\langle p\rangle=0.085$. The curves show the probability that $\nu$ is less than or equal to $x$, where $x$ is any arbitrary value. Results are combined over 100 different network realizations. For each network we select 1 % of nodes as influencers (inferred by one of the heuristics), run the spreading process, track which nodes receive information, and repeat the process $M=10N$ times to account for stochasticity. Red shaded regions denote parts of the distribution where the effective measure is below one, while grey shaded indicate places where the ratio is above one. d, Cumulative distribution of recency for SF networks. e, Fraction of nodes that are worse off with respect to information frequency in $n$ of the seeding heuristics when compared to the benchmark. Error bars are standard deviation over 100 network realizations. f, Fraction of nodes that are worse off with respect to recency.
@@
To formalize our observations from Fig. 1b we first investigate the four influencer maximization heuristics on a testbed of synthetic unweighted and undirected networks with scale-free (SF) degree distributions (see SM Sec. S6 for synthetic networks with normally distributed degree distributions). While perfect SF networks are rarely observed in nature they are powerful simplifications of real-world networks  (?). In order to compare heuristics we construct influencer sets from a fixed finite fraction of the network population—1% of nodes (see SM Sec. S5 for other seed sizes).
@@
Seeding information through random nodes in SF networks results in a near-homogeneous frequency distribution (Fig. 1c black line). In Fig. 1c, a completely equal distribution of information would be characterized by a vertical line in the cumulative probability distribution. (Note that the black line characterizing the random process is not vertical due to the intrinsic variations of network structures.) Using the four influence maximization methods to select influencers, however, results in fundamentally different frequency distributions (Fig. 1c, colored lines). Approximately half of nodes have effective frequency values above one, meaning they receive information more frequently then expected compared to a random process, while the other half receives much less (see also SM Sec. S4). We observe similar results for recency, albeit slightly more polarized (Fig. 1d). Looking across the different influencer heuristics, Fig. 1e, illustrates that if a node is under-informed by one heuristic, it will most likely not be better informed by any other heuristic. On average $42.8\pm 1.7\%$ of nodes are better informed when information is seeded using any influencer maximization heuristics. We say that these nodes are always better off. However, $47.0\pm 2.2\%$ of nodes are consistently left behind, independent of which method is used to select influencers. With respect to recency, Fig. 1f shows an even worse situation, $55.6\pm 2.2\%$ of nodes receive out-of-date information, independent on which information maximization methodology is used.
@@
To understand the implications of information inequalities for real-world networks we look at 10 social networks, encompassing communication, interaction, and collaboration networks varying in size from hundreds to tens of thousands of individuals. The networks are diverse in context, ranging from: face-to-face encounters  (?), connections between households in multiple villages  (?), connections between bloggers and blogs on political topics  (?), digital communication between university students  (?), email communication  (?), online friendships on Facebook  (?), and scientific collaborations  (?, ?) (see SM Sec. S2 for details).
@@
Repeating the analysis from Fig 1 for these real-world networks, we find influencer maximization heuristics to, on average, leave significant portions of the networks in disadvantaged positions in terms of information frequency (Fig. 2a). Overall HD, CHD, and DD result in fairly similar $\nu$ distributions, while selecting influencers according to KC performs worse (with up to 80% of nodes being worse off). This is due to real-world network having a large numbers of cliques, where small discrepancies in shell numbers can result in KC only selecting nodes from a single clique  (?), effectively limiting the diffusion of information. Similar behavior is observed for information recency. Fig. 2b shows that HD, CHD, and DD leave behind a comparable numbers of nodes, while KC consistently performs worse. Summarizing the average reach of influencer heuristics, we find that up to $69.8\%$ of a network might receive information less frequently compared to if it is input at random (Fig. 2c), and information can reach up to $79.2\%$ of individuals slower (Fig. 2d).
@@
There is a connection between access to information and a node’s position in the network. The connection is so strong that a predictive model, based on structural features of as node can accurately predict whether a node will fall into the group of ‘worse off according to all influencer heuristics‘ or ‘better off in all’ (SM Sec. S8). We can, on average, with 97.4% accuracy predict the information status of nodes regarding frequency, and 96.9% for recency. This demonstrates that current influencer heuristics suffer from biases which disadvantage low-connected and peripheral nodes.
@@
Figure 2: Information is unequally distributed in real-world social networks. Here we show results for five of the networks, see SM Fig S9 for results for other networks. a, Cumulative distribution of individual node frequency for networks ordered according to size (number of nodes). Initial seeds contain $1\%$ of network nodes, and results are averaged over $10N$ simulations (see SM Table S1). b, Cumulative distribution of recency for empirical networks. c, Fraction of nodes that are worse off with respect to frequency in $n$ of the seeding heuristics when compared to the random seeding procedure ($\nu^{\text{method}}/\nu^{\text{benchmark}}$). Demonstrating that large parts of social networks are in disadvantaged positions. d, Fraction of nodes that are worse off with respect to recency ($\tau^{\text{method}}/\tau^{\text{benchmark}}$) for $n$ methods.
Fair Influence Maximization
@@
Different strategies can be employed to bridge the information gap. Previous work have shown that instead of using influencer algorithms to identify $s$ individuals, one can select slightly more individuals $s+\epsilon$, but at random  (?). Even for small $\epsilon$-values this has been shown to result in larger cascades. Another solution is to apply acquaintance methods, also called friendship-nomination, which work by selecting a random neighbor of a randomly selected node. These have been used for mass drug administration campaigns  (?), to seed information about maternal and child health  (?), and for inferring centrally located individuals suitable as monitors for detecting large scale disease outbreaks  (?). Lastly, algorithms that use iterative realizations of ICMs to equitably maximize social welfare objective functions have been proposed to bridge the information gap  (?), however, we find them not to be effective at bridging the information gap (see SM Fig. S12).
@@
We embark on a different solution. Traditionally, influence maximization has only focused on maximizing a single objective function, information spread. However, as literature from the field of Artificial Intelligence shows, focusing only on one parameter can lead to troubling and unfair outcomes  (?, ?). Instead, we propose a multi-objective formulation of the fair influence maximization, where both spread and fairness are taken into account in the fitness of a candidate solution (set of selected influencers). We measure fairness as the fraction of nodes that receive information at a higher frequency and speed than what is expected from the benchmark model. We say a node is vulnerable if it receive less information than what can be expected at random ($\nu_{i}^{\text{method}}/\nu_{i}^{\text{benchmark}}<1$). In the following we focus primarily on information frequency, but similar results can be obtained for recency. The more fair an influencer set is, the fewer nodes will be vulnerable, so we maximize the number of non-vulnerable nodes in addition to maximizing spread. Analytically calculating how information will spread from a set of nodes according to the ICM model is a computationally hard problem (NP-hard)  (?). Instead, we use an approximation of the fitness of a influencer set (see SM Sec. S9). To find fairer seeds we then use a genetic algorithm to solve the optimization problem (see Methods).
@@
Fig. 3a,b results for one network, however, our fair influence maximization method works equally well for other real-world networks (Fig. 3e-h, and SM Fig. S15). For other networks (Fig. 3e-h) we find similar results. While influence maximization heuristics identify seed sets which optimize cascade size, we find that these seed sets are not fair in terms of information equality (for numerical results see SM Sec. S10). For all networks it is possible to improve this, with large gains in fairness being achievable with small trade-offs in cascade size. Overall we find that a $1\%$ decrease in cascade size can result in a decrease in number of vulnerable individual by approximately: $1.9\%$ for the network of political blogs, $3.3\%$ for the email communication network, $3.4\%$ for student communication, up to $10.6\%$ for collaboration networks, and $24\%$ for online friendships on Facebook (here we disregard the face-to-face and village networks due to the low number of identified seed sets).
Lastly, for the collaboration networks (Fig. 3h), we find that seed sets identified by influence maximization heuristics are not even close to the Pareto frontier. Meaning they are sub-optimal both in term of cascades sizes and fair information access. In this situation our algorithm can also be used to identify, previously undiscovered, seed sets which optimize cascade-size.
Figure 3: Fair influence maximization for social networks. a, Theoretical Pareto front of optimal influencer sets identified by our multi-objective algorithm for the social network between households in a South Indian village, compared to influence maximization heuristics. We disregard KC as it consistently performs worse, both in terms of fairness and reach compared to the other heuristics. Higher values of non-vulnerable nodes indicate higher values of fairness. b, Numerical evaluation of influencer sets using ICMs. Error bars are given as the standard deviation from 10 realization of $10N$ ICM simulations. c, Edge activations for the set of influencers identified by CHD. Edges are colored and sized according to how often they are activated during $10N$ simulations. Nodes colored black are seed nodes. d, Edge activations for one of the fairer seed sets identified by our algorithm. A comprehensive comparison between the seed sets is available in SM Sec. S10. e-h, Theoretical Pareto fronts for four additional real-world social networks (see SM Fig. S15 for results for remaining five networks, and for numerical results from ICMs).